P(4,2) lies on the curve y=sqrt (x). if Q is (x,sqrt(x)) find the slope of the secant line of PQ for the value of x=3.99. Round to six decimal places.
Answer by stanbon(75887)

P(4,2) lies on the curve y=sqrt (x). if Q is (x,sqrt(x)) find the slope of the secant line of PQ for the value of x=3.99. Round to six decimal places.
----------------------------
You have two points: P(4,2) and Q(3.99,sqrt(3.99))
---------------------------
slope = [sqrt(3.99)-2]/[3.99-4] = -0.002501564/0.01 = -0.2501564...
======================================================================
Cheers,
Stan H.